# Two-Player Zero-Sum Games with Bandit Feedback

### v2

@@ Abstract
We study a two-player zero-sum game in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback. We propose three algorithms based on the Explore-Then-Commit framework. The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration. Our objective is to demonstrate the applicability of ETC in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria. A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret of our proposed algorithms, which has received limited attention in the literature on zero-sum games. Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O(\log(T\Delta^{2})/\Delta)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap. Therefore, our results indicate that ETC-based algorithms perform effectively in adversarial game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
@@ 1 Introduction
Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback. The variant of UCB algorithm for adversarial games under bandit feedback has been analyzed using worst-case regret bounds in [27]. In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game. Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees. More recently, instance-dependent regret bounds have been investigated for the Tsallis-INF algorithm in TPZSGs [18]. However, an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
@@
This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs. While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to adversarial game settings, which makes this the first work to provide its detailed analysis in such a setting. A key motivation for studying ETC is its algorithmic simplicity, which allows for a clear structure.
In particular, we propose three algorithms that adapt the ETC approach to the TPZSG setting and present their instance-dependent regret analyses. The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieves an upper bound of $O(\Delta+\sqrt{T})$ . The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of an $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players. Finally, the last algorithm extends the adaptive elimination approach through non-uniform exploration, which also achieves a regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Consequently, despite the simplicity of the ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings.
@@
Section 2 reviews related work. Section 3 introduces preliminaries and formalizes a two-player zero-sum game setting. In Section 4, we present the ETC in a TPZSG setting and analyze its regret. Section 5 introduces an adaptive elimination algorithm and provides its regret analysis. In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret bound. Section 7 offers empirical results to validate their theoretical performances. Finally, Section 8 discusses our conclusions and future research directions.
@@ 2 Related Work
In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, which makes it essential to estimate the unknown payoffs by previous observations. Various approaches, such as UCB [27] and Tsallis-INF [18], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret. Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments. Additionally, [22] investigates instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
While [27] focuses on the worst-case regret analysis, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium. Compared to [18], our work introduces more refined notions of suboptimality gaps for game instances, whereas their regret analysis focuses solely on the gap between the NE and the action chosen by a player when the opponent plays the equilibrium action. Since their regret bound involves the square of the sum of suboptimality gaps, our instance-dependent analysis gives a tighter bound.
@@ 3.2 Games with Bandit Feedback
The game proceeds in rounds. At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$. The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, with the maximizing player obtaining $r_{t}$ and the minimizing player $r_{t}$. We assume that both players observe each other’s action.
Since the algorithms do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played. We express the action pair played in round $t$ by $(i_{t},j_{t})$. After round $t$, we estimate each entry in the estimated game matrix by computing the average payoff for each action pair $(i,j)$ as
@@
where $n_{ij,t}$ is the number of times the action pair $(i,j)$ played until round $t$, $r_{s}$’s are independently and identically distributed (iid) $\sigma$-subgaussian payoffs. We assume that the estimator is unbiased, which means that $\mathbb{E}[\hat{A}]=A$.
@@ 3.4 Regret Notions
where $i^{t}$ and $j^{t}$ are the actions selected by the row and column player in round $t$, respectively. $R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of the row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret of not choosing the column player’s best action for a given action $i$.
@@
$R_{T}^{\text{max}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}],$
$R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}],$
@@ 4 ETC-TPZSG Algorithm
As a result, the unknown payoff matrix is approximated by the ETC approach, as in [20] for a bandit setting. The ETC algorithm consists of two phases: an exploration phase, in which the player randomly samples actions to gather information with a given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the observations. Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the committing phase the algorithm converges to an optimal action pair which refers to the pure NE.
@@
1:Input:
[…]
3:$m$: number of actions for row player
4:$l$: number of actions for column player
5:$k$: exploration time
6:$T$: time horizon $\triangleright$ $1\leq mlk\leq T$
[…]
8:Initialize: $\hat{A}=[0]^{m\times l}$ $\triangleright$ $\hat{A}$ is estimated game matrix
9:In round t=0,1,2,,mlkt=0,1,2,...,mlk; $\triangleright$ Exploration Phase
10:    Explore each action pair $(i,j)$ in $S_{A}$ $k$ times
11:    Update $\hat{A}$ using (4)
12:In round t=mlk+1,mlk+2,,Tt=mlk+1,mlk+2,...,T; $\triangleright$ Committing Phase
13:    Play an equilibrium $(i^{*},j^{*})$ satisfying the following
14:    property:
@@
The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs, is upper bounded as follows:
@@
where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (16) with the value of the exploration time in (17). Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we find that $\Delta=\frac{4\sigma}{\sqrt{T}}$ is a critical point. If we put it in the regret bound (18), we have the following:
@@ 5 ETC-TPZSG-AE Algorithm
In this section, we extend Algorithm 1 by incorporating adaptive elimination, called ETC-TPZSG-AE. We use the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition, to eliminate action pairs. Specifically, if a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (3) while a $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is challenging.
@@
We set $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$, where $t$ ranges from 0 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$. We have $\hat{\Delta}_{t}\approx\sqrt{(16\sigma^{2}/k_{t})\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}}$ from (21), which implies $\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}\geq 0$. It gives us $\hat{\Delta}_{t}\geq\sqrt{16\sigma^{2}/T}$ and our initialization of $\hat{\Delta}_{t}$ remains consistent with this observation. Then, using $2^{-t+2}\sigma\geq\sqrt{16\sigma^{2}/T}$ and after some calculations, we obtain $t\leq\frac{1}{2}\log_{2}T$. The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise, the number of plays would be zero in the last round.
@@
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big)$ (22)
@@
1:Input:
@@
3:$m$: number of actions for row player
4:$l$: number of actions for column player
5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
[…]
7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
8:In round $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$;
9:    Action Pair Selection:
10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
11:      Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
12:      Update $\hat{A}$ using (4)
13:     Else: $\triangleright$ Committing Phase
14:      Play with a unique action pair in $S_{t}$ until step $T$
15:    Action Pair Elimination:
16:     Remove the action pairs $(i,j)$ NOT satisfying the following for $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}}$ :
@@
17:    Update:
18:     $S_{t+1}$ by "Action Pair Elimination" step
19:     $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
Algorithm 2 ETC-TPZSG-AE
@@
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}+\lambda T\Big).$ (23)
@@
Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$. Specifically, in [18], $\Delta$ refers to the gap between pure NE and the action chosen when the opponent plays the action in equilibrium. Moreover, our regret bound is tighter, as their formulation involves the square of the sum of suboptimality gaps, while ours uses the sum of squares of suboptimality gaps within the logarithmic factor. Moreover, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$, which also allows us to explain why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy.

### v3

@@ Abstract
We study a two-player zero-sum game in which the row player aims to maximize their payoff against an adversarial column player, under an unknown payoff matrix estimated through bandit feedback. We propose three algorithms based on the Explore-Then-Commit (ETC) framework. The first adapts it to zero-sum games, the second incorporates adaptive elimination that leverages the $\varepsilon$-Nash Equilibrium property to efficiently select the optimal action pair, and the third extends the elimination algorithm by employing non-uniform exploration. Our objective is to demonstrate the applicability of ETC in a zero-sum game setting by focusing on learning pure strategy Nash Equilibria. A key contribution of our work is a derivation of instance-dependent upper bounds on the expected regret of our proposed algorithms, which has received limited attention in the literature on zero-sum games. Particularly, after $T$ rounds, we achieve an instance-dependent regret upper bounds of $O(\Delta+\sqrt{T})$ for ETC in zero-sum game setting and $O(\log(T\Delta^{2})/\Delta)$ for the adaptive elimination algorithm and its variant with non-uniform exploration, where $\Delta$ denotes the suboptimality gap. Therefore, our results indicate that ETC-based algorithms perform effectively in zero-sum game settings, achieving regret bounds comparable to existing methods while providing insight through instance-dependent analysis.
@@ 1 Introduction
Several existing studies have advanced our understanding of how players learn to play optimally in zero-sum games with bandit feedback. The variant of UCB algorithm for adversarial games under bandit feedback has been analyzed using worst-case regret bounds in [29]. In particular, worst-case analyses lead to all games as equally hard by ignoring the specific structure of the game. Instance-dependent analysis adapts the regret guarantees to the specific properties of a game to provide more practical performance guarantees. More recently, instance-dependent regret bounds have been investigated in TPZSGs [18, 24]. However, an understanding of instance-dependent regret in TPZSGs under bandit feedback remains incomplete.
@@
This paper investigates the effectiveness of the Explore-Then-Commit (ETC) algorithm and extensions in the context of TPZSGs. While ETC has been widely studied in standard multi-armed bandit (MAB) problems, there has been no prior work specifically analyzing its performance to zero-sum game settings, which makes this the first work to provide its detailed analysis in such a setting. A key motivation for studying ETC is its algorithmic simplicity, which allows for a clear structure. Moreover, our main focus is to show whether ETC and action elimination algorithms is applicable to zero-sum games by offering a refined understanding of regret behavior in TPZSGs under bandit feedback.
In particular, we propose three algorithms that adapt the ETC approach to the TPZSG setting and present their instance-dependent regret analyses. The first one follows the classical ETC framework by uniformly selecting actions for a fixed number of steps and then committing to a pure Nash Equilibria (NE), achieves an upper bound of $O(\Delta+\sqrt{T})$ . The second algorithm incorporates an adaptive elimination (AE), similar to [6], by sequentially eliminating actions that, with high probability, are not part of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), gives us an upper regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Intuitively, this design is expected to converge more rapidly to the optimal strategy by efficiently reducing uncertainty and narrowing the action sets of the players. Finally, the last algorithm extends the adaptive elimination approach through non-uniform exploration, which also achieves a regret bound of $O(\frac{\log(T\Delta^{2})}{\Delta})$. Consequently, despite the simplicity of the ETC framework, our results demonstrate that ETC-based approaches can be highly effective in adversarial settings.
@@
Section 2 reviews related work. Section 3 introduces preliminaries and formalizes a two-player zero-sum game setting. In Section 4, we present the ETC in a TPZSG setting and analyze its regret. Section 5 introduces an adaptive elimination algorithm and provides its regret analysis. In Section 6, we extend the elimination algorithm by using the non-uniform exploration approach and analyze its regret upper bound. Section 7 offers empirical results to validate their theoretical performances. Finally, Section 8 discusses our conclusions and future research directions.
@@ 2 Related Work
In a zero-sum game setting with bandit feedback, players engage in repeated interactions without full knowledge of the payoff matrix, which makes it essential to estimate the unknown payoffs by previous observations. Various approaches, such as UCB [29], Tsallis-INF [18] and a novel algorithm [24], have been proposed to efficiently estimate the unknown payoff matrix while minimizing regret. Learning the game matrix is crucial for converging to equilibrium strategies and ensuring optimal decision-making in adversarial environments. Additionally, [23] and [22] investigate instance-dependent sample complexity bounds for zero-sum games under bandit feedback while [11] examines the convergence rates to NE.
While [29] focuses on the worst-case regret analysis and [24] analyzes instance-dependent regret only for a $2\times 2$ game matrix, [18] provides an instance-dependent regret analysis which achieves an upper bound of $O(\log_{+}(T\Delta^{2})/\Delta)$ in the case of a pure strategy equilibrium. Compared to [18], our work introduces more refined notions of suboptimality gaps for game instances, whereas their regret analysis focuses solely on the gap between the NE and the action chosen by a player when the opponent plays the equilibrium action. Since their regret bound involves the square of the sum of suboptimality gaps, our instance-dependent analysis gives a tighter bound.
@@ 3.2 Games with Bandit Feedback
The game proceeds in rounds. At time $t$, the row player draws an action $i_{t}\in S_{x}$ from a strategy $p_{t}$, and the column player draws $j_{t}\in S_{y}$ from a strategy $q_{t}$. The players then observe a noisy payoff $r_{t}$ with expected value $\mathbb{E}[r_{t}\mid i_{t},j_{t}]=A(i_{t},j_{t})$, where the maximizing player receives $r_{t}$ and the minimizing player receives $r_{t}$. We assume that both players observe each other’s action.
Since the players do not know $A$, we use the notation $\hat{A}(i,j)$ to denote the estimated payoff when the action pair $(i,j)$ is played. We express the action pair played in round $t$ by $(i_{t},j_{t})$. In each round $t$, the estimated game matrix is updated by computing the average payoff for each action pair $(i,j)$ as follows:
@@
where $n_{ij,t}$ is the number of times the action pair $(i,j)$ played until round $t$, $r_{s}$’s are independently and identically distributed (iid) $\sigma$-subgaussian payoffs.
@@ 3.4 Regret Notions
where $i_{t}$ and $j_{t}$ are the actions selected by the row and column player in round $t$, respectively. $R_{T}^{\text{max}}$ represents the expected loss due to not selecting the best possible action of the row player against column player’s action $j$ while $R_{T}^{\text{min}}$ captures the expected regret of not choosing the column player’s best action for a given action $i$.
@@
$R_{T}^{\text{max}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{max}}\mathbb{E}[n_{ij,T}]$
$R_{T}^{\text{min}}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{\text{min}}\mathbb{E}[n_{ij,T}]$
@@ 4 ETC-TPZSG Algorithm
As a result, the unknown payoff matrix is approximated by the ETC approach, as in [20] for a bandit setting. The ETC algorithm consists of two phases: an exploration phase, in which the player randomly samples actions to gather information with a given exploration time $k$, and a commit phase, where the player selects the empirically best action based on the observations. Algorithm 1 presents the implementation of ETC-TPZSG, where each player only observes the payoffs of the action pair they choose during the exploration period, then in the commit phase, the algorithm selects an optimal action pair which refers to the pure NE.
@@
Algorithm 1 ETC-TPZSG
1:Input: $S_{A},m,l,k,\sigma,T(1\leq mlk\leq T)$
2:Initialize: $\hat{A}=[0]^{m\times l}$
3:for $t=1$ to mlkmlk do $\triangleright$ Exploration Phase
4:  Explore each $(i,j)$ in $S_{A}$ $k$ times
5:  Update $\hat{A}$ using (4)
6:end for
7:for $t=mlk+1$ to $T$ do $\triangleright$ Commit Phase
8:  Play an equilibrium $(i^{*},j^{*})$ satisfying the following property:
@@
The Nash regret of Algorithm 1, when interacting with $\sigma$-subgaussian payoffs and $1\leq Nk\leq T$, is upper bounded as follows:
@@
where the first term $T\Delta$ is the worst-case regret and for the other, we combine the regret we obtain in (16) with the value of the exploration time in (17). Then, focusing on $\max\Big\{0,\ln{\frac{\Delta^{2}T}{16\sigma^{2}}}\Big\}$, we obtain $\Delta\geq\frac{4\sigma}{\sqrt{T}}$. Thus, when we put it in the regret bound (18), we have the following:
@@ 5 ETC-TPZSG-AE Algorithm
In this section, we extend Algorithm 1 by incorporating adaptive elimination, called ETC-TPZSG-AE. We use the concept of $\varepsilon$-Nash Equilibrium ($\varepsilon$-NE), which approximately satisfies the standard NE condition, to eliminate action pairs. Specifically, if a pure NE exists at $(i^{*},j^{*})$, then it must satisfy the condition given in (3) while $\varepsilon$-NE satisfies the NE conditions within a tolerance level $\varepsilon$, making it a suitable criterion in learning settings where exact equilibrium identification is challenging.
@@
We set $\hat{\Delta}_{t}=2^{-t+2}\sigma$ in each round $t$, where $t$ ranges from 0 to $\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$, which allows a nonnegative exploration time $k_{t}$ and tolerance level $\varepsilon_{t}$. We have $\hat{\Delta}_{t}\approx\sqrt{(16\sigma^{2}/k_{t})\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}}$ from (21), which implies $\ln{(\hat{\Delta}_{t}^{2}T/16\sigma^{2})}\geq 0$. It gives us $\hat{\Delta}_{t}\geq\sqrt{16\sigma^{2}/T}$ and our initialization of $\hat{\Delta}_{t}$ remains consistent with this observation. Then, using $2^{-t+2}\sigma\geq\sqrt{16\sigma^{2}/T}$ and after some calculations, we obtain $t\leq\frac{1}{2}\log_{2}T$. The division by $e$ is a technical adjustment, which makes the bound safer in the analysis; otherwise, the number of plays would be zero in the last round.
For the property of $\varepsilon$-NE, we define the tolerance parameter as follows:
$\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}}$ (22)
where $\varepsilon_{t}$ reflects how much players are allowed to deviate from the equilibrium and the tolerance over time gradually decreases to ensure convergence towards the NE. Substituting condition $k_{t}\geq\frac{16\sigma^{2}}{\hat{\Delta}_{t}^{2}}\ln{\Big(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}}\Big)}$ into the equation, where $k_{t}$ is defined in (21), we obtain $\varepsilon_{t}\leq\hat{\Delta}_{t}/2$. This bound indicates that as the number of rounds $t$ increases, the tolerance $\varepsilon_{t}$ becomes smaller, forcing the system to narrow the set of action pairs by progressively eliminating those that do not approximate a NE. As a result, over time, the algorithm keeps fewer action pairs, effectively focusing the exploration on action pairs that are close to the optimal one.
@@
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{256\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (23)
@@
Algorithm 2 ETC-TPZSG-AE
1:Input: $S_{A},m,l,\sigma,T(1\leq ml\ll T)$
2:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
3:for $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$ do
4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
5:   Explore each $(i,j)$ in $S_{t}$ $k_{t}$ times, defined in (21)
6:   Update $\hat{A}$ using (4)
@@
8:   Play with a unique action pair in $S_{t}$ until step $T$
9:  end if
10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
@@
11:  $S_{t+1}\leftarrow S_{t}$
12:  $\hat{\Delta}_{t+1}\leftarrow\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
13:end for
@@
$R_{T}\leq\sum_{(i,j)\in S_{A_{1}}}\Big(\Delta_{ij}+\frac{768\sigma^{2}}{\Delta_{ij}}+\frac{256\sigma^{2}}{\Delta_{ij}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\frac{512\sigma^{2}}{\lambda}\Big)+\lambda T.$ (24)
@@
Compared to our regret bound $O(\frac{\log(T\Delta^{2})}{\Delta})$ with the bound $O(\frac{\log_{+}(T\Delta^{2})}{\Delta})$ in [18], although they seem very similar, we observe that our study introduces a more refined notion of the suboptimality gap $\Delta$. Specifically, in [18], $\Delta$ refers to the gap between pure NE and the action chosen when the opponent plays the action in equilibrium. Moreover, our regret bound is tighter, as their formulation involves the square of the sum of suboptimality gaps, while ours uses the sum of squares of suboptimality gaps within the logarithmic factor. Moreover, our elimination algorithm specifically incorporates the extended pure NE property through $\varepsilon$, which also allows us to explain why we achieve a tighter regret bound for a TPZSG setting where both players have a pure strategy.